OpenStudy (anonymous):

what is the slope for x+3y=6 I think it is -1/3 but i just wanna make sure

OpenStudy (anonymous):

Y=mx +b

OpenStudy (anonymous):

Y=1/3+2

OpenStudy (anonymous):

Y=mx +b

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

yeah thats what I thought thanks

OpenStudy (anonymous):

I think reaper deserves a metal